Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 587
Entry time: 12 Feb 2025, 09:53
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 12.02.2025 |
|
|
General Announcement:
- Delivery to instruments restarted this morning with good FEL performance (above the average). HXRSS in SA2 to be done during the day.
- As PRC, Maurizio Vannoni has nothing to report. So far so good.
CTRL:
- Instrument Control Integration team has a retreat today and tomorrow. So far only a "nothing to report" arrived to XO.
DA:
- Nothing to report.
EEE:
- Nothing to report.
VAC:
- Nothing to report; all fine.
XPD:
- Nothing to report.
DET:
- Nothing major;
- Jungfrau MID issue: looks to be permanently solved; further analysis on the reason behind is being done.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- It had to replace a uTCA in SA3 which was causing timing drifts; after replacement the restarted is ongoing and EEE helps. Hope to have delivery today by lunch time.
UNSYS:
- Nothing to report.
SPB:
- Nothing to report.
FXE:
- XTD2 CRL motion issue is solved: Nerea J. informs Leandro Z. did tests with SPB and motor is behaving properly.
- Issue with LPD appeared this morning (no further info on the issue details) but experts are contacted.
MID:
- Nothing to report.
HED:
- Nothing to report.
SCS:
- Thanks for EEE for recent help.
- Lasermet work is done accordingly to LAS.
SQS:
- Not present.
SXP:
- Nothing to report. |